# Msmsfnet: a multi-stream and multi-scale fusion net for edge detection

### v1

@@ II-B Network architecture
II-C Formulation
II-C1 Training phase
Given an image $u$ in the training dataset, notating $\mathbb{G}$ as its associated edge ground truth, where $u=\{u_{j},j=1,\cdots,|u|\}$ denotes the input image and $\mathbb{G}=\{\mathbb{G}_{j},j=1,\cdots,|\mathbb{G}|\}$ denotes the edge ground truth. Notating $\mathbb{W}$ the collection of parameters in all standard network layer (exculding those corresponding to side output layers and the final fusion layer), and notating $\mathbb{W}_{side}=(\mathbb{W}_{side}^{(1)},\mathbb{W}_{side}^{(2)},\mathbb{W}_{side}^{(3)})$ the collection of parameters in those side output layers, the objective function for the side output layers is defined as:
@@ II-C1 Training phase
where $\ell_{side}$ denotes the loss corresponding to the side outputs.
In order to balance the loss between positive/negative classes, a class-balanced cross-entropy loss function as defined in  [26] is used in formula (1):
$\ell_{side}^{(m)}(\mathbb{W},\mathbb{W}_{side}^{(m)})$ $=$
$-\lambda\sum_{j\in\mathbb{G}_{+}}\log\mathbb{P}(\mathbb{G}_{j}=1|u:\mathbb{W},\mathbb{W}_{side}^{(m)})$
$-1.1\times(1-\lambda)\sum_{j\in\mathbb{G}_{-}}\log\mathbb{P}(\mathbb{G}_{j}|u:\mathbb{W},\mathbb{W}_{side}^{(m)}),$ (2)
where $\lambda=\frac{|\mathbb{G}_{-}|}{|\mathbb{G}|}$,and $1-\lambda=\frac{|\mathbb{G}_{+}|}{|\mathbb{G}|}$, $|\mathbb{G}_{-}|$ and $|\mathbb{G}_{+}|$ represent the non-edge and edge label sets. $\mathbb{P}(\mathbb{G}_{j}=1|u:\mathbb{W},\mathbb{W}_{side}^{m})=\sigma(a_{j}^{(m)})\in[0,1]$ with $\sigma(\cdot)$ representing the sigmoid function and $a_{j}^{(m)}$ representing the activation value at pixel $j$. The edge probability map produced by each side layer is thus defined as:
$\hat{\mathbb{G}}_{side}^{(m)}=\sigma(\hat{A}_{side}^{m}),$ (3)
where $\hat{A}_{side}^{(m)}\equiv\{a_{j}^{(m)},j=1,2,\cdots,|\mathbb{G}|$ represent the activations of the output of the side layer.
@@ II-C2 Testing phase
III Experiments
In this section, we demonstrate the efficiency of our proposed msmsfnet by comparing it with several representative CNN-based edge detectors (HED [21], RCF [26], BDCN [32], CATS [33], DexiNed [35]) in three publicly available datasets for edge detection (the BIPEDv2 dataset [35], the BSDS500 dataset [36] and the NYUDv2 dataset) [37]. We show in our experiments that our method achieves better performance than state-of-the-art CNN-based edge detectors in all these datasets when all models are trained from scratch.
III-A Implementation details
@@
The proposed msmsfnet is implemented using the PyTorch library, and trained from scratch using the Adam optimizer [46]. The weights of all convolutional layers are randomly initialized using the Xavier uniform distribution and the biases are all initialized to 0. Weight decay is set to $10^{-12}$. The mini-batch size for training is set to 6. The learning rate is initialized to $10^{-4}$, and decreased by 10 after a fixed number of epochs (10 for the BIPEDv2 dataset, 20 for the BSDS500 dataset, and 20 for the NYUDv2 dataset). The maximum numner of epochs to train is set to 15 for the BIPEDv2 dataset, and 25 for the BSDS500 and NYUDv2 dataset. The optimal performance obtained by each method during training is reported in the following.
III-B Evaluation criteria
@@
The outputs of all CNN-based edge detectors are edge probability maps instead of binary edge maps. Non-maxima suppression, thresholding and edge thining are still required to obtain binary edge maps. Following previous works [21, 26, 32], we adopt the non-maxima suppression step of  [6] to process the outputs of our model. Three criteria are used to compare the performance of different methods: the Optimal Dataset Scale (ODS) F1-score, where a global optimal threshold is chosen to binarize the edge probability maps; the Optimal Image Scale (OIS) F1-score, where an optimal threshold is chosen for each image to binarize the edge probability maps ; and the Average Precision (AP). Higher values of these criteria indicate better performance obtained by a method. Following previous works, precision-recall curves are also ploted to compare the performance of different methods. In the following, we will compare the performance of our msmsfnet with state-of-the-art deep learning based edge detectors using these criteria in three publicly available datasets.
III-C BIPEDv2 dataset
@@
We first compare the performance of our proposed msmsfnet with HED [21], RCF [26], BDCN [32], CATS [33] and DexiNed [35] in the second version of the Barcelona Images for Perceptual Edge Detection (BIPEDv2) dataset [35], as the performance reported by existing algorithms on it does not involve the use of pretrained weights on the ImageNet dataset.
@@
TABLE I: Quantitative comparison of different algorithms on the 50 testing images of the BIPEDv2 dataset.
@@
The BIPEDv2 dataset contains 200 images for training and 50 images for testing. All images are outdoor images and of size $1280\times 720$ ppixels. We use the data augmentation strategy described in  [35] to augment the training dataset, and train all all models from scratch using the augmented training dataset. Quantitative comparisons of different methods on the 50 testing images of the BIPEDv2 dataset are displayed in table I.
From table I we can see that the proposed msmsfnet obtains the highest scores in all three criteria. Minor improvements over existing methods (0.3 percent for ODS F1-score and 0.1 percent for OIS F1-score) are made by our msmsfnet on the test set of BIPEDv2 dataset. The improvents made by our method are relatively larger for the Average Precision (at least 0.9 percent). The precision-recall curves computed by different methods in the test set of BIPEDv2 dataset are displayed in figure 3.
As shown in table I and figure 3, the proposed msmsfnet achieves slightly better performance than existing methods in the BIPEDv2 dataset. This is probably because the edge anotations in the dataset are ’low level edges’, while the multi-representation capability of a model plays a more important role in higher level tasks such as boundary/contour detection. As what we will show in follwing, much larger improvements are obtained by the proposed msmsfnet in the boundary/contour detection datasets. Details about the differences among edges, boundaries, and contours can be found in  [35] (following previous works [21, 26, 25, 32, 33], we do not strictly distinguish between edge and boundary/contour detection and call all these tasks as edge detection).
@@
The edge maps computed by different methods in an image from the test set of BIPEDv2 dataset are displayed in figure 4. From figure 4 we can see that the edge maps computed by HED, RCF and DexiNed are noisy. Many false edges are detected by them. Compared to BDCN and CATS, fewer false edges are detected by our proposed method. Compare to the edge maps computed by other methods, the edge map computed by our method is cleaner and closer to the edge ground truth.
@@ III-D BSDS500 dataset
TABLE II: Quantitative comparisons of different algorithms on the test set of BSDS500 dataset.
@@
msmsfnet 0.816 0.835 0.859
Table II shows the quantitatie comparisons of different methods in the test set of BSDS500 dataset. It is clearly shown that the proposed msmsfnet achieves much higher scores than state-of-the-art edge detectors under the same experimental settings. The ODS F1-score achieved by our methods is at lease 1.3 higher percent than existing methods. In particular, the ODS F1-score computed by our method is 1.3 percent higher than humans, which shows that the perception ability of our model is stronger than humans even without the need to pretrain the model on the ImageNet dataset. The Average Precision computed by our method is at lease 1.8 percent higher than existing methods. The performance gains obtained by our method in the BSDS500 dataset indicate the strong multi-scale representation capability of our model. Precision-recall curves computed by differenet methods are displayed in figure 5. The edge maps computed by different methods in an image from the test set of BSDS500 dataset are shown in figure 6. It is clearly shown that the edge map computed by our method is closer to the edge ground truth. More true edges are detected with fewer false detections. In contrast, the edges detected by BDCN, CATS and DexiNed tend to be fragmented, while many false edges are detected by HED.
Refer to caption Refer to caption
(a) image (b) ground truth
Refer to caption Refer to caption
(c) HED (d) RCF
Refer to caption Refer to caption
(e) BDCN (f) CATS
Refer to caption Refer to caption
(g) DexiNed (h) msmsfnet
@@
III-E NYUDv2 dataset
The NYU-Depth V2 (NYUDv2) dataset [37] contains 1449 pairs of RGB and depth images, which are well aligned, densely labeled and made into the same size. These images have been splitted into three parts: 381 images for training, 414 images for validation and 654 images for testing, The images in the dataset are video sequences of various indoor scenes and recorded by both RGB and Depth camera from Microsoft Kinect. Following previous works [26, 32], both training and validation images are augmented and used to train the models. We adopt the same data augmentation strategy as  [26]. Each model has been trained twice: one trained using RGB images and one trained using HHA features. We compare the performance of different methods using three types of edge predictions produced by each model: one produced by the model trained on RGB images, one produced by the model trained on HHA features, and one produced by averaging the two previous predictions. During evaluation, the maximum tolerance allowed to math the edge predictions with the edge ground truth correctly is increased from 0.0075 to 0.011 as done previous works [21, 26, 32]. Quantitative comparisons of different methods in the test set of NYUDv2 dataset are displayed in table III.
TABLE III: Quantitative comparisons of different algorithms on the test set of NYUDv2 dataset.
@@ III-E NYUDv2 dataset
As shown in table III, our model achieves the highest scores in all experimental settings. When all models are trained using only RGB images, the ODS F1-score computed by our method is at least 1.1 percent higher than those computed by existing methods. The Average Precision computed by our method is at least 2.3 percent higher than those computed by existing methods. When only HHA features are used to train the model, the ODS F1-score computed by our method is at least 1.9 percent higher. The improvements on the Average Precision is much larger, at least 3.5 percent higher than state-of-the-art methods. For the edge predictions computed by combining the information from both RGB images and HHA features, the improvements on the ODS F1-score is relatively minor, only 0.3 percent higher than the best result computed by DexiNed. However, the Average Precision computed by our method is 5.4 percent higher than that of DexiNed. The ODS F1-score computed by our method is at least 1.6 percent higher than the rest methods. Similar improvements have been done on the Average Precision, at least 1.5 percent higher than the rest methods. The precision-recall curves computed by different methods in all experimental settings are displayed in figure 7. The edge maps computed by different methods in an image from the test set of the NYUDv2 dataset can be found in figure 8, figure 9 and figure 10. Figure 8 displays the edge detection results computed by different methods using only RGB images. Figure 9 displays the results computed different methods using only HHA features, while figure 10 displays the results computed using the information from both RGB images and HHA features. It can be seen from these figures that the edges detected by our method are more complete and fewer false edges are detected.
Refer to caption Refer to caption
(a) RGB image (b) ground truth
Refer to caption Refer to caption
(c) HED (d) RCF
Refer to caption Refer to caption
(e) BDCN (f) CATS
Refer to caption Refer to caption
(g) DexiNed (h) msmsfnet
@@
Refer to caption Refer to caption
(a) HHA feature (b) ground truth
Refer to caption Refer to caption
(c) HED (d) RCF
Refer to caption Refer to caption
(e) BDCN (f) CATS
Refer to caption Refer to caption
(g) DexiNed (h) msmsfnet
@@
Refer to caption Refer to caption
(a) RGB image (b) HHA feature
Refer to caption Refer to caption
(c) HED (d) RCF
Refer to caption Refer to caption
(e) BDCN (f) CATS
Refer to caption Refer to caption
(g) DexiNed (h) msmsfnet
Figure 10: Edge detection results computed by different methods using both RGB and HHA features of a test image in the NYUDv2 dataset.
IV Conclusions

### v3

@@ 3.2 Network architecture
3.3 Formulation
3.3.1 Training phase
Given an image $u$ in the training dataset, notating $\mathbb{G}$ as its associated edge ground truth, where $u=\{u_{j},j=1,\cdots,|u|\}$ denotes the input image and $\mathbb{G}=\{\mathbb{G}_{j},j=1,\cdots,|\mathbb{G}|\}$ denotes the edge ground truth. Notating $\mathbb{W}$ the collection of parameters in all standard network layer (excluding those corresponding to side layers and the final fusion layer), and notating $\mathbb{W}_{side}=(\mathbb{W}_{side}^{(1)},\mathbb{W}_{side}^{(2)},\mathbb{W}_{side}^{(3)})$ the collection of parameters in those side layers, the objective function for the side layers is defined as:
@@ 3.3.1 Training phase
where $\ell_{side}^{(m)}$ denotes the loss corresponding to the $m$-th side output.
In order to balance the loss between positive and negative classes, a class-balanced cross-entropy loss function as defined in  [32] is used in formula (1):
$\ell_{side}^{(m)}(\mathbb{W},\mathbb{W}_{side}^{(m)})=-\lambda\sum_{j\in\mathbb{G}_{+}}\log\mathbb{P}(\mathbb{G}_{j}=1|u;\mathbb{W},\mathbb{W}_{side}^{(m)})$
$-1.1\times(1-\lambda)\sum_{j\in\mathbb{G}_{-}}\log\mathbb{P}(\mathbb{G}_{j}=0|u;\mathbb{W},\mathbb{W}_{side}^{(m)}),$ (2)
where $\lambda=\frac{|\mathbb{G}_{-}|}{|\mathbb{G}|}$, and $1-\lambda=\frac{|\mathbb{G}_{+}|}{|\mathbb{G}|}$, $|\mathbb{G}_{-}|$ and $|\mathbb{G}_{+}|$ represent the non-edge and edge label sets. $\mathbb{P}(\mathbb{G}_{j}=1|u;\mathbb{W},\mathbb{W}_{side}^{(m)})=\sigma(a_{j}^{(m)})\in[0,1]$ with $\sigma(\cdot)$ representing the sigmoid function and $a_{j}^{(m)}$ representing the activation value at pixel $j$. The edge probability map produced by each side layer is thus defined as:
$\hat{\mathbb{G}}_{side}^{(m)}=\sigma(\hat{A}_{side}^{(m)}),$ (3)
where $\hat{A}_{side}^{(m)}\equiv\{a_{j}^{(m)},j=1,2,\cdots,|\mathbb{G}|$ represents the activations of the output of the side layer.
@@ 3.3.2 Testing phase
4 Experiments
In this section, we demonstrate the efficiency of our proposed msmsfnet by comparing it with several representative CNN-based edge detectors (HED [28], RCF [32], BDCN [38], CATS [39], DexiNed [25]) in both natural images (three publicly available datasets for edge detection, the BIPEDv2 dataset [25], the BSDS500 dataset [23] and the NYUDv2 dataset [24]) and SAR images (a 1-look SAR dataset simulated using optical images and a 1-look real SAR image). We show in our experiments that our method achieves better performance than state-of-the-art CNN-based edge detectors in all these experiments when all models are trained from scratch. We also show in our experiments that our model is able to achieve competitive performance on the BSDS500 dataset when the pre-trained weights from the ImageNet dataset are used. The proposed method is not compared to UAED  [40] and MUGE  [41] for the following reasons: first, the contributions of UAED  [40] and MUGE  [41] lie mainly on the strategies they proposed to fully exploit the potential of datasets providing multiple annotations for each image and not on the network architecture; second, the strategies they proposed are applicable only when multiple annotations are available for each image and thus not widely useful; third, only one annotation is provided for each image in the BIPEDv2  [25] and NYUDv2  [24] datasets; fourth, since our work focuses on the design of new network architecture for edge detection, the comparison between the performance of different algorithms should be done under the same experimental settings and would be unfair otherwise. Comparisons with the transformer based edge detector EDTER  [42] will be provided in Subsection 4.7.
4.1 Implementation details
@@
The proposed msmsfnet is implemented using the PyTorch library, and trained from scratch using the Adam optimizer. The weights of all convolutional layers are randomly initialized using the Xavier uniform distribution and the biases are all initialized to 0. Weight decay is set to $10^{-12}$. The mini-batch size for training is set to 6. The learning rate is initialized to $10^{-4}$, and decreased by 10 after a fixed number of epochs (10 for the BIPEDv2 dataset, 20 for the BSDS500, NYUDv2 and the simulated SAR dataset). The maximum number of epochs to train is set to 15 for the BIPEDv2 dataset, and 25 for the BSDS500, NYUDv2 and the simulated SAR dataset. The optimal performance obtained by each method during training is reported in the following. The implementation details of the pre-training and fine-tuning experiment on the BSDS500 dataset can be found in Subsection 4.7.
4.2 Evaluation criteria
@@
The outputs of all CNN-based edge detectors are edge probability maps instead of binary edge maps. Non-maxima suppression, thresholding and edge thinning are still required to obtain binary edge maps. Following previous work [28, 32, 38], we adopt the non-maxima suppression step of  [6] to process the outputs of our model. Three criteria are used to compare the performance of different methods: the Optimal Dataset Scale (ODS) F1-score, where a global optimal threshold is chosen to binarize the edge probability maps; the Optimal Image Scale (OIS) F1-score, where an optimal threshold is chosen for each image to binarize the edge probability maps; and the Average Precision (AP). Higher values of these criteria indicate better performance obtained by a method. Following previous work, precision-recall curves are also ploted to compare the performance of different methods. In the following, we will compare the performance of our msmsfnet with state-of-the-art deep learning based edge detectors using these criteria in three publicly available datasets, and a 1-look simulated SAR dataset.
4.3 BIPEDv2 dataset
@@
We first compare the performance of our proposed msmsfnet with HED [28], RCF [32], BDCN [38], CATS [39] and DexiNed [25] in the second version of the Barcelona Images for Perceptual Edge Detection (BIPEDv2) dataset [25], as the performance reported by existing algorithms on it does not involve the use of pre-trained weights on the ImageNet dataset.
@@
Table 1: Quantitative comparison of different algorithms on the 50 testing images of the BIPEDv2 dataset.
@@
The BIPEDv2 dataset contains 200 images for training and 50 images for testing. All images are outdoor images and of size $1280\times 720$ pixels. We use the data augmentation strategy described in  [25] to augment the training dataset, and train all models from scratch using the augmented training dataset. Quantitative comparisons of different methods on the 50 testing images of the BIPEDv2 dataset are displayed in table 1.
From table 1 we can see that the proposed msmsfnet obtains the highest scores in all three criteria. Minor improvements over existing methods (0.3 percent for ODS F1-score and 0.1 percent for OIS F1-score) are made by our msmsfnet on the test set of BIPEDv2 dataset. The improvements made by our method are relatively larger for the Average Precision (at least 0.9 percent). The precision-recall curves computed by different methods in the test set of BIPEDv2 dataset are displayed in figure 3.
@@
As shown in table 1 and figure 3, the proposed msmsfnet achieves slightly better performance than existing methods in the BIPEDv2 dataset. This is probably because the edge anotations in the dataset are ’low level edges’, while the multi-representation capability of a model plays a more important role in higher level tasks such as boundary/contour detection. As what we will show in the following, much larger improvements are obtained by the proposed msmsfnet in the boundary/contour detection datasets. Details about the differences among edges, boundaries, and contours can be found in  [25] (following previous work [28, 32, 31, 38, 39], we do not strictly distinguish between edge and boundary/contour detection and call all these tasks as edge detection).
The edge maps computed by different methods in an image from the test set of BIPEDv2 dataset are displayed in figure 4. From figure 4 we can see that the edge maps computed by HED, RCF and DexiNed are noisy. Many false edges are detected by them. Compared to BDCN and CATS, fewer false edges are detected by our proposed method. Compared to the edge maps computed by other methods, the edge map computed by our method is cleaner and closer to the edge ground truth.
@@ 4.4 BSDS500 dataset
Table 2 shows the quantitative comparisons of different methods in the test set of BSDS500 dataset. It is clearly shown that the proposed msmsfnet achieves much higher scores than state-of-the-art edge detectors under the same experimental settings. The ODS F1-score achieved by our method is at lease 2 percent higher than existing methods. In particular, the ODS F1-score computed by our method is 2 percent higher than humans, which shows that the perception ability of our model is stronger than humans even without the need to pre-train the model on the ImageNet dataset. The Average Precision computed by our method is at least 1.7 percent higher than existing methods. The performance gains obtained by our method in the BSDS500 dataset indicate the strong multi-scale representation capability of our model. Precision-recall curves computed by different methods are displayed in figure 5. The edge maps computed by different methods in an image from the test set of BSDS500 dataset are shown in figure 6. It is clearly shown that the edge map computed by our method is closer to the edge ground truth. More true edges are detected with fewer false detections. In contrast, the edges detected by BDCN, CATS and DexiNed tend to be fragmented, while many false edges are detected by HED.
Table 2: Quantitative comparisons of different algorithms on the test set of BSDS500 dataset.
@@
msmsfnet 0.823 0.840 0.858
Refer to caption Refer to caption Refer to caption Refer to caption
(a) image (b) ground truth (c) HED (d) RCF
Refer to caption Refer to caption Refer to caption Refer to caption
(e) BDCN (f) CATS (g) DexiNed (h) msmsfnet
@@
4.5 NYUDv2 dataset
The NYU-Depth V2 (NYUDv2) dataset [24] contains 1449 pairs of RGB and depth images, which are well aligned, densely labeled and made into the same size. These images have been splitted into three parts: 381 images for training, 414 images for validation and 654 images for testing. The images in the dataset are video sequences of various indoor scenes and recorded by both RGB and Depth camera from Microsoft Kinect. Following previous work [32, 38], both training and validation images are augmented and used to train the models. Each image in the training and validation set is horizontally flipped and rotated to four different angles. Each model has been trained twice: one trained using RGB images and one trained using HHA features. We compare the performance of different methods using three types of edge predictions produced by each model: one produced by the model trained on RGB images, one produced by the model trained on HHA features, and one produced by averaging the two previous predictions. During evaluation, the maximum tolerance allowed to match the edge predictions with the edge ground truth correctly is increased from 0.0075 to 0.011 as done in previous work [28, 32, 38]. Quantitative comparisons of different methods in the test set of NYUDv2 dataset are displayed in table 3.
Table 3: Quantitative comparisons of different algorithms on the test set of NYUDv2 dataset.
@@ 4.5 NYUDv2 dataset
$\text{msmsfnet}_{\text{RGB+HHA}}^{aug+multi}$ 0.763 0.783 0.803
As shown in table 3, our model achieves the highest scores in all experimental settings. When all models are trained using only RGB images, the ODS F1-score computed by our method is at least 1.1 percent higher than those computed by existing methods. The Average Precision computed by our method is at least 2.3 percent higher than those computed by existing methods. When only HHA features are used to train the model, the ODS F1-score computed by our method is at least 1.9 percent higher. The improvements on the Average Precision is much larger, at least 3.5 percent higher than state-of-the-art methods. For the edge predictions computed by combining the information from both RGB images and HHA features, the improvements on the ODS F1-score is relatively minor, only 0.3 percent higher than the best result (computed by DexiNed). However, the Average Precision computed by our method is 5.4 percent higher than that of DexiNed. The ODS F1-score computed by our method is at least 1.6 percent higher than the rest methods. Similar improvements have been done on the Average Precision, at least 1.5 percent higher than the rest methods.
We also show that by using more complicated data augmentation technique and multi-scale testing, the performance of the proposed method (denoted as $\text{msmsfnet}_{*}^{aug+multi}$) can be improved further. Specifically, each image in the training and validation set is horizontally flipped, vertically flipped and flipped in both directions. Then, all images are rescaled to five different scales: $50\%$, $80\%$, $100\%$, $120\%$, and $150\%$. The multi-scale testing technique is the same as the one used in Subsection  4.4. As shown in table  3, significant improvements have been done by the proposed msmsfnet when using more complicated data augmentation technique and multi-scale testing.
The precision-recall curves computed by different methods in all experimental settings are displayed in figure 7. The edge maps computed by different methods in an image from the test set of the NYUDv2 dataset can be found in figure 8 and figure 9. Figure 9 displays the results computed by different methods using only HHA features. It can be seen from these figures that the edges detected by our method are more complete and fewer false edges are detected.
@@
Refer to caption Refer to caption Refer to caption Refer to caption
(a) RGB image (b) ground truth (c) HED (d) RCF
Refer to caption Refer to caption Refer to caption Refer to caption
(e) BDCN (f) CATS (g) DexiNed (h) msmsfnet
@@
Refer to caption Refer to caption Refer to caption Refer to caption
(a) HHA feature (b) ground truth (c) HED (d) RCF
Refer to caption Refer to caption Refer to caption Refer to caption
(e) BDCN (f) CATS (g) DexiNed (h) msmsfnet
@@
4.6 Comparison in SAR images
As described in Section 1, the statistics of SAR images differs greatly from that of natural images. Therefore, fine-tuning the model from the pre-trained weights of the backbone network on the ImageNet dataset is unlikely to improve the edge detection accuracy in real SAR images even in the presence of a real SAR dataset. What’s worse, there is still no real SAR dataset to train CNN models for edge detection. GRHED [9] proposed that we could train CNN models on simulated datasets for edge detection in real SAR images in the absence of a real SAR dataset. It proposed to simulate SAR images by multiplying natural images with 1-look speckle noise. However, strong differences still exist between simulated SAR and real SAR images, making models trained directly using simulated SAR images inefficient for real SAR images [9]. To tackle this issue, GRHED proposed to combine a ratio operator with fully convolutional layers to handle the differences between simulated SAR and real SAR images. This special design, which enables models trained using simulated SAR images to perform efficient edge detection in real SAR images, makes the pre-trained weights of the backbone network on the ImageNet dataset useless because of the ratio operator injected before convolutional layers.
In the following, we demonstrate the efficiency of our proposed msmsfnet by comparing it with a traditional ratio operator GR [55] and state-of-the-art CNN-based edge detectors in both simulated SAR dataset and a 1-look real SAR image. Following the paradigm of GRHED [9], we build GR-HED, GR-RCF, GR-BDCN, GR-CATS, GR-DexiNed, and GR-msmsfnet by combining a ratio operator GR [55] with HED [28], RCF [32], BDCN [38], CATS [39], DexiNed [25] as well as our proposed msmsfnet. We simulate a 1-look SAR dataset with the augmented BSDS500 dataset [23] and the flipped VOC Context dataset [54] as described in  [32] by multiplying the images with 1-look speckle noise [9]. We compare the performance of different methods in the two hundred 1-look images which are simulated from the test set of BSDS500 and a 1-look real SAR image (SLC data of Sentinel-1, ESA, IW mode, date: May 12, 2021, area: Paris, France).
Table 4: Comparison of different methods in the 200 1-look speckled optical images simulated using the test set of BSDS500.
Method ODS OIS AP
GR 0.589 0.615 0.479
GR-HED 0.695 0.711 0.716
GR-RCF 0.693 0.708 0.708
GR-BDCN 0.694 0.710 0.715
GR-CATS 0.695 0.711 0.722
GR-DexiNed 0.697 0.714 0.709
GR-msmsfnet 0.700 0.717 0.724
Quantitative comparisons of different methods in the simulated SAR dataset are shown in Table 4. It is clearly shown in Table 4 that our proposed msmsfnet achieves higher scores in all criteria, which demonstrates the efficiency of our proposed msmsfnet. Besides, the scores computed by the deep learning based edge detectors are significantly higher (around 10 percent higher in terms of ODS F1-score and OIS F1-score, and more than 20 percent higher in terms of AP) than those of GR [55], which demonstrates the benefits of exploiting convolutional neural networks to improve the performance of edge detectors in images corrupted by strong noise.
Refer to caption Refer to caption Refer to caption Refer to caption
(a) Sentinel-1 image (b) GR (c) GR-HED (d) GR-RCC
Refer to caption Refer to caption Refer to caption Refer to caption
(e) GR-BDCN (f) GR-CATS (g) GR-DexiNed (h) GR-msmsfnet
Figure 10: Edge detection results computed by different edge detection algorithms in a 1-look real SAR image (Sentinel-1) of size $512\times 512$ pixels.
We further compare the performance of different algorithms in a 1-look real SAR image of size $512\times 512$ pixels. The edge maps computed by different algorithms are displayed in Figure 10. It is clearly shown in Figure 10 that our proposed method GR-msmsfnet achieves the best edge detection results with more true edges detected and much fewer false detections. As shown in Figure 10-(h), the edges detected by our proposed GR-msmsfnet are more complete. Compared to the traditional ratio operator GR, GR-CATS also achieves better edge detection result with more complete edges and fewer false detections. However, the edge detection results computed by GR-HED, GR-RCF, GR-BDCN and GR-DexiNed in the real SAR image are very poor and worse than that computed by GR, although the scores computed by them in the simulated SAR dataset are significantly higher than those of GR. The efficiency of GR-msmsfnet and GR-CATS in the real SAR image shows that the strategy of combining a traditional ratio operator with fully convolutional neural networks successfully enables the models trained using simulated SAR images to work in real SAR images. In contrast, the poor edge detection results computed by GR-HED, GR-RCF, GR-BDCN and GR-DexiNed in the real SAR image poses an important issue of CNNs: its generalization capability. This is indeed worthy of study in future work.
4.7 Pre-training and fine-tuning
In order to demonstrate the efficiency of the proposed method, we show that our method is able to achieve competitive performance in the BSDS500 dataset when the pre-trained weights from the ImageNet dataset are used to initialize model’s parameters. To obtain the pre-trained weights on the ImageNet dataset, we modify out network architecture for image classification. We add a max-pooling layer, a global average pooling layer and a 1000-way fully connected layer after the last msmsfblock so that it can perform the task of image classification on the ImageNet dataset. All the side layers are removed. The modified model is trained using the training dataset of the ImageNet-1k dataset, which contains around 1.28 million images. All training Images (RGB) are randomly resized and then cropped into size of $224\times 224$ pixels to be used as the input of the network. Randomly horizontal flipping is used for data augmentation. We train our model using the SGD optimizer and use the cross-entropy loss as the loss function. During training, the learning rate is initially set to 0.01, and divided by 10 every 60 epochs. The training batch size is set to 256, momentum is set to 0.9, and the weight decay is set to 0.0001. Our model reaches its best performance on the validation set of the ImageNet dataset after training for 145 epochs. It achieves $70.91\%$ top-1 accuracy, which is slightly lower (less than one percent) than the VGG net and around five percent lower than the ResNet-50 net. It should be noted that our aim here is just to obtain the pre-trained weights from the ImageNet dataset which will be used to fine-tune our model for edge detection. The performance of our model may be improved if a better training strategy is used.
After obtaining the pre-trained weights, our model is fine-tuned for edge detection on the the BSDS500 dataset. During fine-tuning, we find it beneficial to adopt the fusion strategy of BDCN. Thus, we modify our network architecture following the fusion strategy of BDCN to obtain better performance. During fine-tuning, our experimental settings are the same as  [32, 38]. The training dataset and the loss function is the same as those used in Subsection 4.4. Following previous work  [28, 32, 38], we use SGD optimizer to fine-tune our model for edge detection. The learning rate is initially set to $10^{-6}$, and divided by 10 after 12 epochs. Momentum is set to 0.9 and weight decay is set to 0.0002. We train our model for 20 epochs and the best performance (msmsfnetpre-training) is reported in table  5. The performance obtained by msmsfnet with random initialization (msmsfnetrandom) is also reported.
We compare the performance of our model with BDCN, which reported the best performance on the BSDS500 dataset among state-of-the-art CNN-based edge detectors, and EDTER, a transformer-based edge detector which reports the best performance for edge detection on the BSDS500 dataset.
[…]
For EDTER, we report three results. One is copied from the paper (EDTERpaper), one is obtained through our re-training of EDTER following the training settings in the paper (EDTERre-training), and the third one is obtained by training EDTER from scratch (EDTERrandom). In order to get the best performance achievable by EDTER with random initialization, we train EDTER using both Adam optimizer and SGD optimizer. In both training settings, the model is trained for 30 epochs, and the learning rate is divided by 10 after 20 epochs. the weight decay is set to $10^{-12}$. The mini-batch size for training is 6. Three initial learning rates are used for the Adam optimizer, namely $10^{-3}$, $10^{-4}$, and $10^{-5}$. Five initial learning rates are used for the SGD optimizer: $10^{-3}$, $10^{-4}$, $10^{-5}$, $10^{-6}$, and $10^{-7}$. The best performance obtained by EDTER are reported (see table 5).
Table 5: Quantitative comparison between BDCN, EDTER and the proposed msmsfnet on the test set of BSDS500 dataset. The pre-trained weights for BDCN and msmsfnet are obtained by pre-training the backbone network on the ImageNet-1k dataset, while the pre-trained weights for EDTER are obtained from the ImageNet-22k dataset, namely the ImageNet-21k dataset plut the ImageNet-1k dataset.
Methods Pre-training ODS OIS AP
[…]
EDTERpaper ImageNet-22k 0.848 0.865 0.903
EDTERre-training ImageNet-22k 0.847 0.863 0.855
msmsfnetpre-training ImageNet-1k 0.839 0.853 0.875
[…]
EDTERrandom - 0.736 0.769 0.776
msmsfnetrandom - 0.823 0.840 0.858
The performance obtained by BDCN, EDTER, and the proposed msmsfnet in the BSDS500 dataset can be found in table  5. From table 5 we can see that the proposed msmsfnet achieves superior performance than BDCN when the pre-trained weights are used to initialize the models’ parameters. Even though EDTER achieves the best performance on the BSDS500 dataset in all three criteria, the pre-trained weights which are used for parameter initialization are obtained by pre-training the backbone network on the ImageNet-22k dataset, namely the ImageNet-21k dataset plus the ImageNet-1k dataset. The ImageNet-21k dataset contains around 14 million images and is much larger than the ImageNet-1k dataset, which contains around 1.28 million images. Therefore, the comparison between the performance of EDTER and the proposed msmsfnet when using the pre-trained weights is not fair, as much more data has been used to train EDTER. When all models are trained from scratch, it is clearly shown in table 5 that the proposed msmsfnet achieves much better performance than BDCN and EDTER. In particular, the performance achieved by EDTER is far below than that of BDCN and msmsfnet, around 7 percent lower than BDCN and 9 percent lower than msmsfnet in terms of ODS F1-score. This can be expected as it has been proven in  [43] that the vision transformer requires sufficient amount of data to obtain better performance than CNNs. In the absence of sufficient amount of data for training, the CNN-based models are better choices as they can obtain better performance.
5 Conclusions